The area of a gymnasium floor is 264 square yards. The floor is 11 yards wide. How long is the floor?
a)253 yards
b)24 yards
c)11 yards
d) 242 yards

Answer by Zenock(15)

If the area is A then the A=L*W so L=A/W. Note that all measurements need to be in the proper units. Here everything is in yards so just plug in 264 for A and 11 for W.
Example: Some area A (like a gymnasium floor) is 300 square yards. The area is 15 yards wide (W). How long (L) is the area.
We use the equation L=A/W, so L=300/15=20 yards wide.
This should give you all the information you need to solve the problem.
If you need more help, please post again.